# Small Drafts, Big Verdict: Information-Intensive Visual Reasoning via Speculation

### v1

@@ 1 Introduction
Recent advances in large vision-language models (VLMs) have delivered impressive performance on tasks such as image captioning and general visual question answering (VQA) (Li et al., 2025; Fu et al., 2024). However, these models encounter challenges in information-intensive images that densely interleave diverse textual annotations (legends, labels, captions) with fine-grained graphical elements (charts, diagrams, plots) across multiple scales and formats (Su et al., 2025b). Addressing this task requires two interdependent capabilities (Figure 1; Ke et al., 2025): (i) comprehensive and precise localization, which involves not only pinpointing the exact positions of critical cues in densely populated layouts but also ensuring that all query-relevant regions are identified; (ii) multi-hop reasoning, which chains visual analysis—encompassing colors, shapes, and spatial relationships—with textual evidence, thereby integrating dispersed cues into a coherent and complete answer. As each reasoning step builds on the accuracy of the previous one, any intermediate error can propagate through the entire chain, making the overall process highly error-sensitive and difficult to correct retrospectively.
@@
Existing work tackles information-intensive visual reasoning with search‐based zoom‐in pipelines that enlarge local regions for detailed reasoning. Specifically, learning-based methods train reinforcement learning policies to guide zoom operations iteratively (Zheng et al., 2025; Su et al., 2025a; Fan et al., 2025; Zhang et al., 2025b). Enhancing its performance would demand costly fine-grained supervision. Moreover, training-free methods perform cropping based on internal attention or confidence scores (Zhang et al., 2025a; Shen et al., 2024; Wang et al., 2025c). Yet in dense layouts, we find these signals correlate weakly with true relevance, misleading the model into visually similar but irrelevant areas. Consequently, these tool-driven designs fail to capture all evidence for multi-hop reasoning, leaving the core challenges of information-intensive visual reasoning unsolved.
@@
To overcome these limitations, we propose Speculative Verdict (SV), a training-free framework inspired by speculative decoding that combines small draft visual experts with a large verdict model (Leviathan et al., 2023). The framework operates in two stages (Figure 2): (1) Draft stage: multiple lightweight VLMs serve as draft experts, each generating a reasoning path that offers diverse localization candidates; (2) Verdict stage: a large VLM acts as a strong verdict, which receives the reasoning paths as contextual evidence, distinguishes the correct information, and outputs the final answer. SV directly tackles core challenges through complementary strengths: draft experts expand evidence coverage across scattered regions, while the verdict prevents error propagation by synthesizing these multiple perspectives. Importantly, unlike using a large proprietary model to reason over every image section, SV invokes the verdict only once to yield a concise final answer, thereby minimizing computational cost while effectively recovering correct answers. To further balance accuracy and efficiency, SV introduces a consensus expert selection mechanism in the draft stage, ensuring that only reasoning paths with strong agreement are forwarded to the verdict.
@@
We evaluate Speculative Verdict on information-intensive VQA benchmarks, including InfographicVQA (Mathew et al., 2021), ChartMuseum (Tang et al., 2025), and ChartQAPro (Masry et al., 2025), which demand reasoning over dense textual and visual content. As a training-free framework, SV consistently outperforms strong open-source models, large proprietary models, and perception-focused search methods while remaining cost-efficient. In particular, SV yields average gains of 4% over small VLMs as draft experts and 10% over GPT-4o (Hurst et al., 2024) as verdict. Beyond overall gains, SV successfully corrects 47-53% of cases where majority voting or the verdict model alone fails, thereby reducing vulnerability to error propagation in information-intensive visual reasoning. Furthermore, SV surpasses all baselines on HR-Bench 4K (Wang et al., 2025b), a benchmark for high-resolution visual perception, underscoring its effectiveness in challenging multimodal reasoning scenarios.
@@
Figure 1: Examples of correct reasoning paths for information-intensive image VQA tasks. They illustrate distinct paths: (a) focuses on the localization of a specific chart, symbol identification, and complementary reasoning from a single percentage value; (b) focuses on keyword-based localization, evidence aggregation from multiple entries across the entire image, and cross-entity sorting to select the minimum.
@@ 2 Related Work
Vision-Language Model Reasoning with Tools. Recent research has explored enhancing VLM perception by manipulating input images with zooming operations to locate relevant regions (Hu et al., 2024). (1) Prompting-based methods exploit internal signals of VLMs to decide where to zoom. ViCrop (Zhang et al., 2025a) leverages models’ attention maps to highlight query-related regions, thereby generating automatic visual crops. Other works perform tree-based search, where models evaluate candidate sub-images with confidence scores to iteratively narrow down to relevant regions (Shen et al., 2024; Wang et al., 2025c). However, such signals align poorly with the required evidence in information-intensive images, since queries often require reasoning across multiple dispersed regions. (2) Reinforcement learning approaches instead optimize policies that interleave visual zooming with textual reasoning (Zheng et al., 2025; Su et al., 2025a; Fan et al., 2025; Zhang et al., 2025b). By calling zooming tools within the agentic framework, these methods adaptively crop regions and concatenate them into the reasoning trajectory, enabling more active evidence gathering. Yet these methods still fall short on information-intensive images, requiring costly task-specific training to scale.
@@
Speculative Decoding. Speculative decoding is a draft-then-verify decoding paradigm to accelerate LLM inference (Xia et al., 2024). Specifically, it utilizes a draft model to generate future tokens, and a larger target model verifies them via parallel rejection sampling. Beyond the vanilla setting, recent work extends acceptance from token-level equivalence to step-level semantic similarity to speed up reasoning (Yang et al., 2025; Pan et al., 2025; Fu et al., 2025b; Liao et al., 2025). Collaborative decoding via Speculation (Fu et al., 2025a) further applies speculative decoding with multiple draft LLMs by verifying proposals against a combined distribution of the drafts and the target, yielding greater speedups than standard ensembling. However, these adaptations primarily target speed in LLM inference and also do not address the challenges of vision-language reasoning.
Large Language Model Ensemble. Majority voting aggregates answers by frequency, but fails when the correct solution is produced by a minority. Universal Self-Consistency (Chen et al., 2023) mitigates this failure mode by prompting the LLM to select the most consistent candidate across samples. Further, learned aggregators read multiple rationales and synthesize them to recover minority-correct information (Qi et al., 2025; Zhao et al., 2025). However, these approaches focus on text-only ensembling. In vision-language reasoning, supervision of ensembling is not cost-effective since multimodal complexity requires costly, fine-grained annotations.
@@ 3 Speculative Verdict
Speculative decoding is an inference‐time optimization originally developed to mitigate the latency of autoregressive generation (Leviathan et al., 2023). The approach employs a draft-then-verify paradigm: (i) a small, fast draft model proposes one or more future tokens speculatively, and (ii) a large, accurate base model verifies these proposals in parallel, accepts or revises the proposals, and generates output that is consistent with the base model’s distribution  (Xia et al., 2024; Zhang et al., 2024). This token-level process speeds up inference by committing several tokens at once, while maintaining quality by discarding continuations that diverge from the base model’s distribution.
@@ 3.2 Draft Stage
Chain-of-Thought (CoT) prompting exposes models’ intermediate reasoning steps in an explicit, stepwise form (Wei et al., 2022). This is critical for information-intensive VQA, where solving a question requires a sequence of localization, evidence extraction, and analytic operations (Figure 1). However, current VLMs often lack fine-grained perception and localization on densely annotated images, and existing tool-driven zoom-in methods are ineffective as elaborated in Section 2. We therefore utilize multiple VLMs to produce reasoning paths rather than a single direct answer, so that the subsequent verdict can verify and synthesize structured evidence. Concretely, given an image-question pair $(x,q)$, we select $m$ lightweight VLMs $\{M_{1},\dots,M_{m}\}$ as draft experts from a pool of $k$ candidate VLMs via a consensus-based selection mechanism (detailed in Section 3.4). Each selected expert $M_{i}$ is then prompted with a CoT template to output a reasoning path $r_{i}$.
@@ 3.4 Consensus Expert Selection
Consensus Expert Selection Strategy. We adopt a cross-all strategy that selects the $m$ VLMs with the strongest consensus, measured by the lowest consensus scores, from the pool of $k$ candidates. As described in Section 3.2, these $m$ selected VLMs then become the draft experts to generate detailed reasoning paths forwarded to the verdict (Figure 3 illustrates this process). By aggregating agreement across all peers, this strategy provides a holistic measure of reliability. It thus yields a subset of reasoning paths that are well-grounded and compact in size, balancing informativeness and efficiency.
@@ 4.1 Setups
Configuration Details. We set the draft pool size to $k=5$ considering efficiency and select $m=3$ draft experts in our main experiments. Ablation studies over different $m$ values are reported in Section 4.4. The draft pool consists of the following VLMs for expert selection: Qwen2.5-VL-7B-Instruct (Bai et al., 2025), MiMo-VL-7B-RL (Xiaomi, 2025), InternVL3-8B (Zhu et al., 2025), GLM-4.1V-9B-Thinking (Team et al., 2025b), Ovis2.5-9B (Lu et al., 2025). These models are chosen as candidate VLMs based on their strong performance on multimodal benchmarks and their diverse architectural designs. For the verdict models, we employ GPT-4o (Hurst et al., 2024) and Qwen2.5-VL-72B-Instruct respectively, given their superior ability in visual reasoning. In particular, for information-intensive image benchmarks, we preprocess images with PP-StructureV3 (Cui et al., 2025) to produce a layout-preserving structured format (see Appendix E.2 for details), provided together with the original image as auxiliary input to the verdict model.
@@
Baselines. We compare SV with proprietary models GPT-4o and GPT-4o-mini, and the large open-source model Qwen2.5-VL-72B-Instruct as it is one of our verdicts. We also evaluate SV against draft experts mentioned above. These baselines are evaluated under the same chain-of-thought prompting template in Appendix H. Additionally, we include DeepEyes (Zheng et al., 2025) as a representative tool-driven baseline with zoom-in operations.
@@
Benchmarks. We evaluate SV on three information-intensive benchmarks and extend the evaluation to a representative high-resolution benchmark, providing a comprehensive assessment of fine-grained visual reasoning: InfographicVQA (Mathew et al., 2021), ChartMuseum (Tang et al., 2025), ChartQAPro (Masry et al., 2025) and HR-Bench 4K (Wang et al., 2025b). InfographicVQA collects infographics with an average high resolution over 2k, designed to test reasoning over layout, graphical and textual content, including operations such as counting, sorting, and basic arithmetic. ChartMuseum and ChartQAPro introduce substantially greater visual reasoning complexity by covering a broad spectrum of real-world chart types and question formats, revealing a large performance gap between current Large VLMs and humans. These benchmarks require models to visually ground relevant regions, extract information, and conduct reasoning to answer queries.
@@
We further assess generalization to high-resolution images on HR-Bench 4K. It comprises two sub-tasks: FSP (Fine-grained Single-instance Perception) and FCP (Fine-grained Cross-instance Perception), stressing small-object perception and cross-instance reasoning under high-resolution inputs.
@@
Table 1: Results on test sets of four benchmarks. InfographicVQA, ChartMuseum, and ChartQAPro are information-intensive VQA benchmarks, while HR-Bench 4K focuses on high-resolution perception. We compare SV against closed-source VLMs, open-source VLMs, and the tool-driven method, with all results reproduced by ourselves. The best results for each benchmark are highlighted in bold and the second-best results are underlined.
@@
Figure 4: SV’s correction ability on verdict’s error cases across information-intensive benchmarks (GPT-4o as verdict). We consider only cases where the verdict itself fails, to isolate SV’s independent correction capacity. For each benchmark, three bars denote expert correctness categories (majority-correct, minority-correct, and zero-correct), defined by how many selected experts provide the correct answer. Within each category, the bars are split into the proportion corrected by SV (dark) versus not corrected (light). More details can be found in Appendix C.
@@ 4.2 Results on Information-Intensive Benchmarks
(i) SV shows consistent gains over all strong draft experts’ baselines, with improvements of 3.6% on InfographicVQA, 1.3% on ChartMuseum, and 6.6% on ChartQAPro with GPT-4o as verdict. SV also achieves comparable gains with Qwen2.5-VL-72B-Instruct as a verdict.
(ii) Importantly, SV enables strong error correction beyond simple answer aggregation. Figure 4 analyzes SV’s performance on cases where the verdict itself fails, categorized by expert correctness (minority-correct, majority-correct, zero-correct). Across benchmarks, SV recovers 47-53% of minority-correct cases, where few draft experts are correct and the verdict alone also fails (case in Figure 3). Moreover, SV even recovers 2.5-4.5% of zero-correct cases, where neither the drafts nor the verdict answers correctly (case in Appendix G). In these cases, SV succeeds because errors in information-intensive visual reasoning are often decomposable, enabling SV to extract partially correct components from different draft reasoning paths while rejecting misleading cues. Thus, SV achieves effective correction where traditional ensemble methods fail.
@@
(iv) SV substantially outperforms representative tool-driven pipeline DeepEyes, with gains of +12.9% on InfographicVQA, +21.3% on ChartMuseum, and +11.3% on ChartQAPro. This gap arises because DeepEyes is strong in local grounding but less effective when reasoning over dense textual and visual content. For example, it often focuses on text spans or legends rather than full regions needed for analytical operations, and its zoom-in calls are sometimes redundant or misdirected (see Appendix F for error analysis). As a result, it struggles with global comparison and dispersed evidence synthesis. In contrast, SV’s reasoning-path synthesis enables it to integrate evidence across regions reliably without relying on predefined tool-based visual search.

### v2

@@ 1 Introduction
Recent advances in large vision-language models (VLMs) have delivered impressive performance on tasks such as image captioning and general visual question answering (VQA) (li2025survey; fu2024mme). However, these models encounter challenges in information-intensive images that densely interleave diverse textual annotations (legends, labels, captions) with fine-grained graphical elements (charts, diagrams, plots) across multiple scales and formats (su2025thinking). Addressing this task requires two interdependent capabilities (Figure 1; explain): (i) comprehensive and precise localization, which involves not only pinpointing the exact positions of critical cues in densely populated layouts but also ensuring that all query-relevant regions are identified; (ii) multi-hop reasoning, which chains visual analysis—encompassing colors, shapes, and spatial relationships—with textual evidence, thereby integrating dispersed cues into a coherent and complete answer. As each reasoning step builds on the accuracy of the previous one, any intermediate error can propagate through the entire chain, making the overall process highly error-sensitive and difficult to correct retrospectively.
@@
Existing work tackles information-intensive visual reasoning with search‐based zoom‐in pipelines that enlarge local regions for detailed reasoning. Specifically, learning-based methods train reinforcement learning policies to guide zoom operations iteratively (zheng2025deepeyes; su2025pixel; fan2025grit; zhang2025chain). Enhancing its performance would demand costly fine-grained supervision. Moreover, training-free methods perform cropping based on internal attention or confidence scores (zhang2025mllms; shen2024zoomeye; rap). Yet in dense layouts, we find these signals correlate weakly with true relevance, misleading the model into visually similar but irrelevant areas. Consequently, these tool-driven designs fail to capture all evidence for multi-hop reasoning, leaving the core challenges of information-intensive visual reasoning unsolved.
@@
To overcome these limitations, we propose Speculative Verdict (SV), a training-free framework inspired by speculative decoding that combines small draft visual experts with a large verdict model (leviathan2023fast). The framework operates in two stages (Figure 2): (1) Draft stage: multiple lightweight VLMs serve as draft experts, each generating a reasoning path that offers diverse localization candidates; (2) Verdict stage: a large VLM acts as a strong verdict, which receives the reasoning paths as contextual evidence, distinguishes the correct information, and outputs the final answer. SV directly tackles core challenges through complementary strengths: draft experts expand evidence coverage across scattered regions, while the verdict prevents error propagation by synthesizing these multiple perspectives. Importantly, unlike using a large proprietary model to reason over every image section, SV invokes the verdict only once to yield a concise final answer, thereby minimizing computational cost while effectively recovering correct answers. To further balance accuracy and efficiency, SV introduces a consensus expert selection mechanism in the draft stage, ensuring that only reasoning paths with strong agreement are forwarded to the verdict.
@@
We evaluate SV on information-intensive VQA benchmarks, including InfographicVQA (mathew2021infographicvqa), ChartMuseum (tang2025chartmuseum), and ChartQAPro (masry2025chartqapro), which demand reasoning over dense textual and visual content. As a training-free framework, SV consistently outperforms strong open-source models, large proprietary models, and perception-focused search methods while remaining cost-efficient. In particular, SV yields average gains of 4% over small VLMs as draft experts and 10% over GPT-4o (hurst2024gpt) as verdict. Beyond overall gains, SV successfully corrects 47-53% of cases where majority voting or the verdict model alone fails, thereby reducing vulnerability to error propagation in information-intensive visual reasoning. Furthermore, SV surpasses all baselines on HR-Bench 4K (wang2025divide), a benchmark for high-resolution visual perception, underscoring its effectiveness in challenging multimodal reasoning scenarios.
@@
Figure 1: Examples of correct reasoning paths for information-intensive VQA tasks. They illustrate distinct paths: (a) focuses on the localization of a specific chart, symbol identification, and reasoning from a single percentage value; (b) focuses on keyword-based localization, evidence aggregation from multiple entries across the entire image, and cross-entity sorting to select the minimum.
@@ 2 Related Work
Vision-Language Model Reasoning with Tools. Recent research has explored enhancing VLM perception by manipulating input images with zooming operations to locate relevant regions. (1) Prompting-based methods exploit internal signals of VLMs to decide where to zoom. ViCrop (zhang2025mllms) leverages models’ attention maps to highlight query-related regions, thereby generating automatic visual crops. Other works perform tree-based search, where models evaluate candidate sub-images with confidence scores to iteratively narrow down to relevant regions (shen2024zoomeye; rap). However, such signals align poorly with the required evidence in information-intensive images, since queries often require reasoning across multiple dispersed regions. (2) Reinforcement learning approaches instead optimize policies that interleave visual zooming with textual reasoning (zheng2025deepeyes; su2025pixel; fan2025grit; zhang2025chain). By calling zooming tools within the agentic framework, these methods adaptively crop regions and concatenate them into the reasoning trajectory, enabling more active evidence gathering. Yet these methods still fall short on information-intensive images, requiring costly task-specific training to scale.
General Vision-Language Model Reasoning. Recent work has also explored other paradigms that enhance VLM reasoning. Prompt-enhanced VLMs use chain-of-thought prompting to articulate intermediate observations and sub-goals, yielding more structured reasoning (llavacot; mitra2024compositional; visualcot). RL-enhanced methods further optimize these trajectories via supervised fine-tuning and reinforcement learning, inspired by reasoning-oriented models such as GPT-o1 and DeepSeek-R1 (visionr1; deepperception; o1; deepseekr1). Recently, agentic frameworks treat the VLM as a planner that decomposes queries and actively chooses actions, either invoking explicit visual tools or performing implicit latent-space reasoning (wu2024v; hu2024visual; cogcom; machine; foresight). However, they remain vulnerable to imprecise or weakly supervised visual operations and error propagation.
LMM-as-a-Judge. Large multimodal models (LMMs) increasingly serve as general-purpose evaluators for vision-language tasks (gptjudge; ge2025mllm; li2025vl). Specifically, LMM judges are prompted or trained to score candidates, produce rankings, or select the best answer given the task context, and instruction (llavacritic). These judges can deliver fine-grained evaluations for open-ended generation and reasoning tasks, and are increasingly used as scalable supervision signals for stages such as alignment, retrieval, and reasoning (li2025generation). In our framework, the verdict model acts as an off-the-shelf multimodal judge that filters informative cues from diverse drafts and synthesizes an answer on information-intensive images.
@@
Speculative Decoding. Speculative decoding is a draft-then-verify decoding paradigm to accelerate LLM inference (xia2024unlocking). Specifically, it uses a draft model to generate future tokens, and a larger target model verifies them via parallel rejection sampling. Recent work extends acceptance from token-level equivalence to step-level semantic similarity to speed up reasoning (yang2025speculative; pan2025specreason; fu2025scaling; liao2025reward). Collaborative decoding via Speculation (fu2025fast) further applies speculative decoding with multiple draft LLMs by verifying proposals against a combined distribution of drafts and target, yielding greater speedups. However, they mainly target speed in LLM inference and do not address visual reasoning challenges.
@@ 3 Speculative Verdict
Speculative decoding is an inference‐time optimization originally developed to mitigate the latency of autoregressive generation (leviathan2023fast). The approach employs a draft-then-verify paradigm: (i) a small, fast draft model proposes one or more future tokens speculatively, and (ii) a large, accurate base model verifies these proposals in parallel, accepts or revises the proposals, and generates output that is consistent with the base model’s distribution  (xia2024unlocking; zhang2024beyond). This token-level process speeds up inference by committing several tokens at once, while maintaining quality by discarding continuations that diverge from the base model’s distribution.
@@ 3.2 Draft Stage
Chain-of-Thought (CoT) prompting exposes models’ intermediate reasoning steps in an explicit, stepwise form (wei2022chain). This is critical for information-intensive VQA, where solving a question requires a sequence of localization, evidence extraction, and analytic operations (Figure 1). However, current VLMs often lack fine-grained perception and localization on densely annotated images, and existing tool-driven zoom-in methods are ineffective as elaborated in Section 2. We therefore utilize multiple VLMs to produce reasoning paths rather than a single direct answer, so that the subsequent verdict can verify and synthesize structured evidence. Concretely, given an image-question pair $(x,q)$, we select $m$ lightweight VLMs $\{M_{1},\dots,M_{m}\}$ as draft experts from a pool of $k$ candidate VLMs via a consensus-based selection mechanism (detailed in Section 3.4). Each selected expert $M_{i}$ is then prompted with a CoT template to output a reasoning path $r_{i}$.
@@ 3.4 Consensus Expert Selection
Consensus Expert Selection Strategy. We adopt a cross-all strategy that selects the $m$ VLMs with the strongest consensus, measured by the lowest consensus scores, from the pool of $k$ candidates. As described in Section 3.2, the $m$ selected VLMs become the draft experts to generate detailed reasoning paths forwarded to the verdict (Figure 3 illustrates this process). By aggregating agreement across all peers, this strategy provides a holistic measure of reliability. It thus yields a subset of reasoning paths that are well-grounded and compact in size, balancing informativeness and efficiency.
@@ 4.1 Setups
Configuration Details. We set the draft pool size to $k=5$ considering efficiency and select $m=3$ draft experts in our main experiments. Ablation studies over different $m$ values are reported in Section 4.5. The draft pool consists of the following VLMs for expert selection: Qwen2.5-VL-7B-Instruct (bai2025qwen2), MiMo-VL-7B-RL (Yue2025MiMo), InternVL3-8B (zhu2025internvl3), GLM-4.1V-9B-Thinking (glm), Ovis2.5-9B (lu2025ovis2). These models are chosen as candidate VLMs based on their strong performance on multimodal benchmarks and their diverse architectural designs. For the verdict models, we employ GPT-4o (hurst2024gpt) and Qwen2.5-VL-72B-Instruct respectively, given their superior ability in visual reasoning. In particular, for information-intensive image benchmarks, we preprocess images with PP-StructureV3 (cui2025paddleocr) to produce a layout-preserving structured format, provided together with the original image as auxiliary input to the verdict model.
@@
Baselines. We compare SV with proprietary models GPT-4o and GPT-4o-mini, and the large open-source model Qwen2.5-VL-72B-Instruct as it is one of our verdicts. We also evaluate SV against draft experts mentioned above. These baselines are evaluated under the same chain-of-thought prompting template in Appendix M. Additionally, we include DeepEyes (zheng2025deepeyes) and Pixel-Reasoner (su2025pixel) as representative tool-driven baselines with zoom-in operations.
@@
Benchmarks. We evaluate SV on three information-intensive benchmarks and extend the evaluation to a representative high-resolution benchmark, providing a comprehensive assessment of fine-grained visual reasoning: InfographicVQA (mathew2021infographicvqa), ChartMuseum (tang2025chartmuseum), ChartQAPro (masry2025chartqapro) and HR-Bench 4K (wang2025divide). InfographicVQA collects infographics with an average high resolution over 2k, designed to test reasoning over layout, graphical and textual content, including operations such as counting, sorting, and basic arithmetic. ChartMuseum and ChartQAPro introduce substantially greater visual reasoning complexity by covering a broad spectrum of real-world chart types and question formats, revealing a large performance gap between current Large VLMs and humans. These benchmarks require models to visually ground relevant regions, extract information, and conduct reasoning to answer queries.
@@
We further assess generalization to high-resolution perception on HR-Bench 4K. It comprises two sub-tasks: FSP (Fine-grained Single-instance Perception) and FCP (Fine-grained Cross-instance Perception), stressing small-object perception and cross-instance reasoning. We also test on two additional multimodal reasoning benchmarks, TallyQA (acharya2019tallyqa) and MathVista (lu2023mathvista), covering open-ended counting with relational reasoning and mathematical visual reasoning.
@@
Table 1: Results on test sets of four benchmarks. InfographicVQA, ChartMuseum, and ChartQAPro are information-intensive VQA benchmarks, while HR-Bench 4K focuses on high-resolution perception. We compare SV against closed-source, open-source VLMs, and tool-driven methods. $\dagger$ denotes results reported in the original papers and all other results are reproduced by ourselves. The best results for each benchmark are highlighted in bold and the second-best results are underlined.
@@
  $\Delta$ (vs. Qwen2.5-VL-72B-Instruct) +2.5 +7.5 +2.3 +2.5
Table 2: Results on additional multimodal reasoning benchmarks. We evaluate on MathVista-testmini and 1000 randomly sampled complex questions from TallyQA to assess generalization.
Model Param Size TallyQA-Complex Acc MathVista Acc
Closed-source VLMs
GPT-4o 75.4 65.1
Open-source VLMs
Qwen2.5-VL-Instruct 7B 72.4 68.2
MiMO-VL-RL 7B 72.0 80.3
InternVL3 8B 70.1 72.7
GLM-4.1V-Thinking 9B 73.9 78.9
Ovis2.5 9B 71.9 77.3
SV w/ GPT-4o Verdict 76.9 82.9
  $\Delta$ (vs. GPT-4o) +1.5 +17.8
Figure 4: SV’s correction ability on verdict’s error cases across information-intensive benchmarks (GPT-4o as verdict). We consider only cases where the verdict itself fails, to isolate SV’s independent correction capacity. For each benchmark, three bars denote expert correctness categories (majority-correct, minority-correct, and zero-correct), defined by how many selected experts provide the correct answer. Within each category, the bars are split into the proportion corrected by SV (dark) versus not corrected (light). More details can be found in Appendix D.
@@ 4.2 Results on Information-Intensive Benchmarks
(i) SV shows consistent gains over all strong draft experts’ baselines, with improvements of 3.6% on InfographicVQA, 1.3% on ChartMuseum, and 6.7% on ChartQAPro with GPT-4o as verdict. SV also achieves comparable gains with Qwen2.5-VL-72B-Instruct as a verdict.
(ii) Importantly, SV enables strong error correction beyond simple answer aggregation. Figure 4 analyzes SV’s performance on cases where the verdict itself fails, categorized by expert correctness (minority-correct, majority-correct, zero-correct). Across benchmarks, SV recovers 47-53% of minority-correct cases, where few draft experts are correct and the verdict alone also fails (case in Figure 3). Moreover, SV even recovers 2.5-4.5% of zero-correct cases, where neither the drafts nor the verdict answers correctly (case in Appendix L). In these hard cases, SV exploits complementary reasoning strengths across draft experts (e.g., extraction, localization, color matching) (see detailed analysis in Appendix I). This complementarity allows the verdict to synthesize partially correct evidence across reasoning paths while rejecting misleading signals. Thus, SV achieves effective correction where traditional ensemble methods fail.
@@
(iv) SV substantially outperforms representative tool-driven pipeline DeepEyes and Pixel-Reasoner. Specifically, it improves over DeepEyes by +12.9% on InfographicVQA, +21.3% on ChartMuseum, and +11.3% on ChartQAPro and it also exceeds Pixel-Reasoner by clear margins. While these methods benefit from zoom-in operations, such operations are often under-triggered or misdirected on dense infographics (see Appendix K). Thus, it struggles with global comparison and dispersed evidence synthesis. Yet, SV’s reasoning-path synthesis enables it to integrate evidence across regions reliably without relying on predefined tool-based visual search.